A scale model of a library is 5 inches long and 3 inches wide. The actual length of the library is 135 feet. What is the width?
Found 2 solutions by mathstutor458, nyc_function:
Answer by mathstutor458(57)

A scale model of a library is 5 inches long and 3 inches wide.
The actual length of the library is 135 feet.
which means,
5inches = 135 feet
1 inch =135/5 = 27 feet
3 inches wide = 3 * 27 = 81 feet width --->answer

Answer by nyc_function(2741)

Let w = width

5/135 = 3/w

135 * 3 = 5w

405 = 5w

405/5 = w

81 = w

The width is 81 feet.